OpenStudy (anonymous):

URGENT! Exam, need to calculate lim x^x as x approaches 0+

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0^{+}} x ^{x}\]

OpenStudy (anonymous):

2nd question) Need to calculate rate of change of length of square whose area equal to 100m2, where rate of change \[\frac{ dA }{ dt }=1 \frac{ m ^{2} }{ s }\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0^{+}} x ^{x}=L\] \[\lim_{x \rightarrow 0^{+}} \ln(x ^{x})=\ln(L)\] \[\lim_{x \rightarrow 0^{+}} x\ln(x)=\ln(L)\]

OpenStudy (anonymous):

\[ x\ln(x) = \frac{\ln(x)}{x^{-1}} \]

OpenStudy (anonymous):

You can use l'Hospitals at that point.

OpenStudy (anonymous):

combinig with implicit differentiation?

OpenStudy (anonymous):

?? I don't know what you mean.

OpenStudy (anonymous):

that equals to 0?

OpenStudy (anonymous):

Yes, but that gives you the equation \[ 0=\ln(L) \]And you really want to find \(L\) remember?

OpenStudy (anonymous):

so our L is equal to 1?

OpenStudy (anonymous):

Yeah!

OpenStudy (anonymous):

great thx)

OpenStudy (anonymous):

what about second question?))

OpenStudy (anonymous):

Close this question and ask your next one in a new quesiton.

OpenStudy (anonymous):

ok